FLUORESCENCE SPECTRA AND LIFETIMES OF THE NH_{2} ^{2}A_{1}- ^{2}B_{1} SYSTEM: EXPERIMENT AND THEORY

C. PETRONGOLO, Dipartimento di Chimica, Universita di Siena, Via A. Moro 2, I-53100 Siena, Italy; H. FAN, I. IONESCU, D. KUFFEL, AND \underline{S. A. REID}, Department of Chemistry, Marquette University, Milwaukee, WI 53233.

We report new calculations and measurements of the fluorescence properties of the title system, which is important in astrochemical processes. Calculations were done omitting the molecular rotation and employing recent ab initio potential energy surfaces. Dispersed fluorescence spectra show extensive X^{2}B_{1} vibrational progressions that depend on the initial A^{2}A_{1} state. Observed and calculated (0,ν′_{2},0) lifetimes are in good accord, save for ν′_{2}=4 (bent molecule notation), and calculated (1, ν′_{2}-2,0) lifetimes are longer than the (0,ν′_{2},0) ones. The electronic transition moment determined from a fit of the observed and calculated lifetimes is in good agreement with ab initio predictions, and the calculated laser-induced fluorescence spectrum will be compared with experimental absorption data.